OpenStudy (anonymous):

NEED HELP! f(x) =3x+2; g(x)=3x-5 find f/g.

OpenStudy (anonymous):

if f(x) =3x +1 g(x) = 3x -4 \[\frac{ f(x) }{ g(x) } = \frac{ 3x + 2 }{ 3x -5 }\]

OpenStudy (anonymous):

can you do the domain of that problem please?

OpenStudy (anonymous):

3x + 5 =. 0 X = 3/5 3* 3/5 + 2 9/5 + 2 18/10 + 20/10 = 38/10 19/5

OpenStudy (anonymous):

k thank you I got it